OpenStudy (anonymous):

2(x+4) = 30 A. 7 B. 11 C. 15 D. 19

OpenStudy (thesecret20111):

2(x + 4) = 30 x + 4 = 15 x = 11

OpenStudy (anonymous):

\[ 2(x+4)=30\\ x+4=15\\ x=11 \]